OpenStudy (anonymous):

Find the solution of the inequality. |x-5|greater than or equal to 3

OpenStudy (jamesj):

So geometrically, it's all numbers x which are a distance of 3 or more from 5. What numbers are those?

OpenStudy (anonymous):

no its this sign > with a line under it so thats why i put greater than or equal to

OpenStudy (jamesj):

Yes, so at a distance 3 or more from 5. For example 8 and all numbers greater than 8. Hence one part of the solution is the real numbers \( x \) such that \( x \geq 8 \). Now, what's the other part?

OpenStudy (anonymous):

I have no clue I don't understand this whole question at all

OpenStudy (jamesj):

Let's step back. What is the definition of |x|?

OpenStudy (anonymous):

?

OpenStudy (jamesj):

What does |x| mean?

OpenStudy (anonymous):

um again i don't know

OpenStudy (jamesj):

Ok. |x| is called the absolute value of x. It measures how far the number x is from zero on the real number line. How far is 1 from 0? 1. Hence |1| = 1. How far is -1 from 0? 1. Hence |-1| = 1. How far is 0 from 0? 0. Hence |0| = 0. Now what is |2| and |-2| ?

OpenStudy (anonymous):

0?

OpenStudy (jamesj):

No, 2 is a distance of 2 units away from zero. Hence |2| = 2. What then is |-2| = ?

OpenStudy (anonymous):

2

OpenStudy (jamesj):

Yes. Now what is |-5|?

OpenStudy (anonymous):

5

OpenStudy (jamesj):

Yes. And what is |2-1|? This is the distance between 2 and 1

OpenStudy (anonymous):

1?

OpenStudy (jamesj):

Yes. And what is the distance between 1 and -1? |1-(-1)| = ... ?

OpenStudy (anonymous):

0?

OpenStudy (jamesj):

Is 1 = -1? I don't think so. Try again. |1-(-1)| = ... ?

OpenStudy (anonymous):

1

OpenStudy (jamesj):

Draw a number lilne. Locate -1 and locate +1. How far apart are they?

OpenStudy (anonymous):

well its -1 then 0 then 1

OpenStudy (jamesj):

Yes, so they are a distance of 2 apart. I.e., |1-(-1)| = |1+1| = 2.

OpenStudy (anonymous):

ohhh ok

OpenStudy (jamesj):

Now, what number or numbers x are a distance 2 apart from 1? i.e., |x-1| = 2.

OpenStudy (anonymous):

1?

OpenStudy (jamesj):

How far apart is 1 from 1? A distance of zero. Hence if x=1, then |x-1| = 0. We want numbers x such that |x-1| = 2. What's one number x that is a distance 2 away from 1?

OpenStudy (jamesj):

look at a number line.

OpenStudy (anonymous):

ughhh i just don't understand this

OpenStudy (jamesj):

Draw a number line. Mark 0 and 1 and the numbers -5 to 5 on it. Now look at 1. What are the numbers a distance 2 to the left of 1 and 2 to the right of 1?

OpenStudy (anonymous):

x is less than or equal to 8 or x is greater than or equal to 2 this is what I put on my test

OpenStudy (jamesj):

We're coming back to your problem next. But answer this one first. If x is a real number with the property that |x-1| = 2 i.e., x is a distance of 2 away from +1, what values can x have?

OpenStudy (anonymous):

-1 and 3?

OpenStudy (jamesj):

3 is one such number. 3 is a distance 2 away from +1. This makes sense as |3-1| = |2| = 2. Now, what's the other number x such that |x-1| = 2 ?

OpenStudy (jamesj):

Yes -1 and 3.

OpenStudy (jamesj):

returning now to your problem, we are told that \[ |x-5| \geq 3 \] Let's first solve \[ | x - 5| = 3 \] What two numbers satisfy this last equation?

OpenStudy (jamesj):

i.e., x is a distance 3 from +5.

OpenStudy (anonymous):

so that would be 2?

OpenStudy (jamesj):

That's one number. What's the other?

OpenStudy (anonymous):

-2?

OpenStudy (jamesj):

No. the difference between -2 and 5 is 7. I.e., |(-2)-5| = |-7| = 7. Try again. |x-5| = 3

OpenStudy (anonymous):

is it 8

OpenStudy (jamesj):

Yes of course. 8 and 2 are both a distance 3 from 5. i.e., if |x-5| = 3, then x = 3 or 8.

OpenStudy (jamesj):

Now you want real numbers x such that \[ |x-5| \geq 3 \] So the distance is equal to or \( \it greater \) than 3. What's a number more than 3 away from 5?

OpenStudy (anonymous):

4

OpenStudy (jamesj):

9 is one such number. So is 10 and 11 and 20 and 200 and 2000 and 50000 and 100000. So is 8.5 and 8.1 and 8.00001 and 8.0000000001.

OpenStudy (jamesj):

Hence *every* number greater than 8 is more than a distance of 3 away from 5.

OpenStudy (anonymous):

so 9?

OpenStudy (jamesj):

Look at the number line. *EVERY* number greater than 8 is more than a distance 3 away from 5. Hence EVERY number greater than or equal to 8 satisfies \[ |x-5| \geq 3 \] make sense?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

im working on it might be getting there but slowly lol

OpenStudy (jamesj):

Now, look on the other side of 5. What are the numbers x which are a distance of 3 or more from 5?

OpenStudy (jamesj):

We know that 2 is one of them. What other numbers?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

is that a question or statement?

OpenStudy (jamesj):

Draw a number line. Mark on it 0 and 5 and 10 and -5 and -10.

OpenStudy (jamesj):

Tell me when you've done that.

OpenStudy (jamesj):

In fact, mark on it all the numbers from -10 to 10.

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

ok i got it

OpenStudy (jamesj):

ok. Now look at 5. What numbers are 3 or more away from 5. Well, on the right hand side, clearly 8 is at least a distance 3 away from 5. And in fact, so must every number greater than 8. So far so good?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Now, to the left of 5, what number is *exactly* a distance of 3 away from 5?

OpenStudy (anonymous):

2?

OpenStudy (jamesj):

Yes 2. Now staying to the left of 5, what numbers are a distance of more than 3 away from 5?

OpenStudy (anonymous):

uhh.. going 3 distance away from would be 2 or if your asking to go 3 more away from five would be -1

OpenStudy (jamesj):

I am asking for numbers to the left of 5 on the number line, in other words for numbers less than 5, what is an example of a number which is a distance of more than 3 away from 5?

OpenStudy (jamesj):

For example, 1 is at a distance of 4 from 5. Hence x=1 satisfies \[ |x-5| \geq 3 \] Yes?

OpenStudy (anonymous):

so it would be 4?

OpenStudy (jamesj):

because in fact |1-5| = |-4| = 4. That is, the distance between 5 and 1 is 4. Now what are all the numbers less than 5 which are at a distance of more than 3? Look at the number line. You know that 2 is *exactly* 3 from 5. What numbers are more than a distance of 3 from 5?

OpenStudy (anonymous):

3 & 4?

OpenStudy (anonymous):

omg im so confused

OpenStudy (jamesj):

the distance between 3 and 5 is 2: |3-5| = |-2| = 2 The distance between 4 and 5 is 1: |4-5| = |-1| = 1 So no, that cannot be right.

OpenStudy (jamesj):

Again, what numbers are *exactly* a distance of 3 away from 5? 8 and 2. Agreed?

OpenStudy (anonymous):

yes, it is and would it be 1?

OpenStudy (jamesj):

So if the question was: for what values of x is |x-5| = 3 the answers would be 2 and 8, and we'd be finished. But that's not the problem.

OpenStudy (jamesj):

The problem is: For what values of x is \[ |x-5| \geq 3 \] That is, what numbers x are at least a distance of 3 away from 5.

OpenStudy (anonymous):

sorry im taking all of your time im just dumb i guess..

OpenStudy (jamesj):

Now, clearly 8 is at least 3 away from 5 and hence every number greater than 8 is more than a distance of 3 away from 5. So far so good?

OpenStudy (anonymous):

yea

OpenStudy (jamesj):

Now what other number is *exactly* a distance of 3 away from 5?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

i dont get how there is another number exactly a distance of 3 away from 5

OpenStudy (jamesj):

Yes 2. Now what other numbers to the left of 5 are more than 3 units away from 5?

OpenStudy (jamesj):

Look at the number line. You have to go 3 units from 5 to 8. And 3 units from 5 to 2. Absolute value measures distance, not direction, not plus/minus. All we care about here is measuring distance.

OpenStudy (jamesj):

I don't know what that means. But +1 is more than 3 units away from 5. So is 0, so is -1. So is 1.5, so is 1.9. In fact, every number less than 2 is more than 3 units distance from 5. Make sense?

OpenStudy (anonymous):

thats more less what i just said every number away from 2 is more than 3 units away from 5

OpenStudy (jamesj):

less than 2, right. So let's pull this altogether: What are all the real numbers x such that \[ |x-5| \geq 3 \]?

OpenStudy (anonymous):

2 and 8

OpenStudy (jamesj):

No, that's the solution to the question: What are the real numbers x such that \[ |x-5| = 3 \] ?

OpenStudy (anonymous):

thats what the question is about finding the solution

OpenStudy (jamesj):

Your original question was: "Find the solution of the inequality. |x-5|greater than or equal to 3" So you need the inequalities as well. In other words, you don't want just the x such that \[ |x-5| = 3 \] We also want the x such that \[ |x-5| > 3 \]

OpenStudy (jamesj):

In other words, we don't just want the x which are exactly a distance 3 from 5, we also want all the x which are a distance greater than 3 from 5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be |2-5|=3

OpenStudy (jamesj):

x=2 is one solution. But it's not all the solutions.

OpenStudy (jamesj):

Look at your number line.

OpenStudy (jamesj):

We want all of the numbers x which are a distance of 3 or more from 5.